# Capacity Bounds for Broadcast Channels with Bidirectional Conferencing Decoders

### v2

@@ Abstract
The two-user broadcast channel (BC) with receivers connected by bidirectional cooperation links of finite capacities, known as conferencing decoders, is considered. A novel capacity region outer bound is established based on multiple applications of the Csiszár-Körner identity. Achievable rate regions are derived by using Marton’s coding as the transmission scheme, together with different combinations of decode-and-forward and quantize-bin-and-forward strategies at the receivers. It is shown that the outer bound coincides with the achievable rate region for a new class of semi-deterministic BCs with degraded message sets; for this class of channels, one-round cooperation is sufficient to achieve the capacity. Capacity result is also derived for a class of more capable semi-deterministic BCs with both common and private messages and one-sided conferencing. For the Gaussian BC with conferencing decoders, if the noises at the decoders are perfectly correlated (i.e., the correlation is either 1 or -1), the new outer bound yields exact capacity region for two cases: i) BC with degraded message sets; ii) BC with one-sided conferencing from the weaker receiver to the stronger receiver. When the noises have arbitrary correlation, the outer bound is shown to be within half a bit from the capacity region for these same two cases. Finally, for the general Gaussian BC, a one-sided cooperation scheme based on decode-and-forward from the stronger receiver to the weaker receiver is shown to achieve the capacity region to within $\frac{1}{2}\log(\frac{2}{1-|\lambda|})$ bits, where $\lambda$ is the noise correlation. An interesting implication of these results is that for a Gaussian BC with perfectly negatively correlated noises and conferencing decoders with finite cooperation link capacities, it is possible to achieve a strictly positive rate using only infinitesimal amount of transmit power.
@@
Broadcast channel, bidirectional cooperation, conferencing decoders, capacity region, deterministic channel, Gaussian broadcast channel
@@ I-A Motivation
The situation is completely different if the receiver noises are correlated. Specifically, if the receiver noises are perfectly positively correlated, then cooperation would gain nothing, because $Y_{1}=Y_{2}$ in this case; as each receiver already knows what the other receiver knows, exchanging information does not help. However, if the receiver noises are perfectly negatively correlated, then cooperation can help significantly. Consider the case where $Y_{1}=X+Z$ and $Y_{2}=X-Z$, and the conferencing links have infinite capacities $C_{12}=C_{21}=\infty$. Then, the BC with conferencing decoders would have infinite capacity at any positive transmit power $P$, because having access to both $Y_{1}$ and $Y_{2}$ allows the noises to be cancelled completely.
What happens if the noises are perfectly negatively correlated, but the conferencing links have finite capacity, e.g., $C_{12}=C_{21}=1$ bit per channel use? The analysis becomes more delicate and this is one of the main topics of this paper. It turns out that an interesting phenomenon emerges—it is possible to achieve a strictly positive rate (i.e., 1 bit in the example above), even with an infinitesimally small transmit power. In contrast to the single-user Gaussian channel where $P\to 0$ implies $C\to 0$, for this Gaussian BC with perfectly negatively correlated noises and conferencing decoders of fixed cooperative link capacities, we can have $P\to 0$, while $C\to 1$. In other words, the energy efficiency for transmitting one bit of information can be infinite!
Figure 1: The capacity of a Gaussian channel (left) goes to zero if the transmit power $P$ goes to zero at a fixed noise variance $N$. In a Gaussian BC with perfectly negatively correlated noises at the receiver (right), conferencing links of fixed capacities $C_{12}$ and $C_{21}$ between the receivers can allow a strictly positive rate to be achieved even at an infinitesimally small transmit power.
@@ I-C Main Contributions
We then propose achievability strategies for the BC with conferencing decoders. In [20, Theorem 2] an achievable region is derived for the BC with bidirectional cooperation by applying Marton’s coding at the transmitter and the compress-and-forward cooperative scheme at both receivers. Another achievable region is given in [21] based on coding strategies for the partially cooperative relay broadcast channels. A third achievable region is given in [23, Appendix B] for the BC with one-sided cooperation between receivers, which is derived by applying Marton’s coding at the transmitter and decode-and-forward as the cooperative protocol. All of these achievable schemes are however insufficient to either derive new capacity results for the discrete memoryless BCs, or approximate capacity results for the Gaussian BC.
@@ II Channel Model
For the BC with conferencing decoders, a length-$n$ code with $\Gamma$ conferencing rounds is described as follows. The transmitter encodes independent messages $M_{0}$, $M_{1}$, and $M_{2}$, which are uniformly distributed over the sets $[1:2^{nR_{0}}]$, $[1:2^{nR_{1}}]$, and $[1:2^{nR_{2}}]$, respectively, into a codeword and sends over the channel according to the following:
@@ III-A Converse
for some joint PDFs $P(u,v,x)$, where $U,V\rightarrow X\rightarrow Y_{1},Y_{2}$ forms a Markov chain and $|\mathcal{U}|\leq|\mathcal{X}|+2$ and $|\mathcal{V}|\leq|\mathcal{X}|+2$. The set $\mathcal{R}_{o}$ constitutes an outer bound on the capacity region.
@@ III-B Achievability
$\zeta_{1}$ =$=\{\alpha_{1}C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2})\}^{+}$ (17)
$\zeta_{2}$ =$=\{C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})\}^{+}$ (18)
for some joint PDFs $P(u,v,w,x)P(\hat{y}_{1}|u,w,y_{1})P(\hat{y}_{2}|y_{2})$ and $\alpha_{1}\in[0,1]$. The convex closure of the set $\mathcal{R}_{i}^{(1)}$ is achievable.
@@
The achievable rate region $\mathcal{R}_{i}^{(1)}$ given in Theorem 2 clearly includes both of the regions previously presented in [20, Theorem 2] and [23, Appendix B] as subsets. Moreover, after some algebraic computation, one can show that $\mathcal{R}_{i}^{(1)}$, restricted over the distributions of the form $P(u,v,w,x)P(\hat{y}_{1}|w,y_{1})P(\hat{y}_{2}|y_{2})$, includes the achievable rate region of [21, Theorem 3] as a subset. Therefore, $\mathcal{R}_{i}^{(1)}$ encompasses these previously known results.
@@
Next, we present a second achievable rate region based on a combination of decode-and-forward and quantize-bin-and-forward at one receiver first and then quantize-bin-and-forward at the other receiver as cooperation strategy. This strategy may be advantageous when part of the message is already decodable at the first receiver before conferencing.
@@
$\eta_{1}$ =$=\{C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2})\}^{+}$ (40)
$\eta_{2}$ =$=\{\alpha_{2}C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})\}^{+}$ (41)
for some joint PDFs $P(u,v,w,x)P(\hat{y}_{1}|u,w,y_{1})P(\hat{y}_{2}|w,y_{2})$ and $\alpha_{2}\in[0,1]$. The convex closure of the set $\mathcal{R}_{i}^{(2)}$ is achievable.
@@
The cooperation strategy to achieve the rate region $\mathcal{R}_{i}^{(2)}$ is to apply Marton’s coding with a combination of decode-and-forward and quantize-bin-and-forward at the receiver $Y_{2}$ first and then quantize-bin-and-forward at the receiver $Y_{1}$. At the receiver $Y_{2}$, the capacity of the conferencing link $C_{21}$ is divided into two parts. A fraction $\bar{\alpha}_{2}C_{21}$ is devoted for decode-and-forward and the remaining $\alpha_{2}C_{21}$ is devoted for quantize-bin-and-forward scheme. A complete proof of Proposition 2 can be found in Appendix C. ∎
@@
One can obtain an alternative achievable rate region by exchanging the cooperative protocol at the receivers in the acheivability scheme of Proposition 2, i.e., to apply a combination of decode-and-forward and quantize-bin-and-forward at the receiver $Y_{1}$ first and then quantize-bin-and-forward at the receiver $Y_{2}$.
@@ III-C Capacity Regions for Specific Channels
The achievability is derived by setting $U=X$, $W=V$, and $\hat{Y}_{2}=Y_{2}$ in $\mathcal{R}_{i}^{(1)}$. The converse proof is readily given by $\mathcal{R}_{o}$. ∎
@@
Figure 3: Capacity region and cut-set outer bound for the BC in Example 2, where $Z$ is a binary noise with ${P(Z=0)=0.2}$.
Unlike the previous example, where the capacity region coincides with the cut-set bound, the capacity region for this BC is characterized using an auxiliary variable $V$. In Fig. 3, we plot the capacity region along with the cut-set outer bound for the channel in Example 2 with $P(Z=0)=0.2$ for two cases: $(C_{12},C_{21})=(0,0.9)$ and $(C_{12},C_{21})=(0.1,0.9)$, by searching over all possible $P(v,x)$. As can be seen, the capacity region lies strictly below the cut-set bound in this example. Moreover, $C_{12}$ strictly improves the capacity region, although $Y_{1}$ only observes the noise. In fact, the constraint (69) (which arises from (60)) plays a crucial role in characterizing the capacity region. Without (69), the remaining constraints simply reduce to the cut-set bound.
For this channel, the optimal coding strategy is as follows. Note that the receiver $Y_{1}$ (which is supposed to decode both common and private messages) does not receive any information from the transmitter directly. The optimal cooperation strategy is that the receiver $Y_{2}$ first sends a compressed version of its received signal to the receiver $Y_{1}$ through the digital link $C_{21}$. Next, the receiver $Y_{1}$ decodes both messages using the information received (and its own signal which is in fact the channel noise) and then forwards part of the common message (using the variable $V$) to the receiver $Y_{2}$ through the link $C_{12}$. Lastly, the receiver $Y_{2}$ decodes the common message using its received signal.
It is worthwhile to make the following observation. For Example 2, as the receiver $Y_{1}$ observes the channel noise only, one might think that a cooperative scheme in which $Y_{1}$ applies quantize-bin-and-forward and $Y_{2}$ applies decode-and-forward is the right strategy. However, it turns out that such a scheme is not optimal when the message to $Y_{2}$ is a degraded version of the message to $Y_{1}$. This is in contrast to Example 1, in which the message set degrades in the opposite direction and performing quantize-bin-and-forward at the receiver that observes the noise only is the capacity-achieving strategy.
As far as we know, the above results are the first cases where a combination of quantize-bin-and-forward and decode-and-forward strategies yields an optimal bidirectional cooperation protocol. There is no previously known capacity result in the literature for channels with bidirectional cooperation between the receiver. Moreover, these results demonstrate that a one-round cooperation scheme is sufficient for achieving capacity for this class of channels.
@@
The achievability is derived by setting $U=X$, $W=V$, and $\hat{Y}_{2}=Y_{2}$ in $\mathcal{R}_{i}^{(1)}$.
For the converse, note that due to the Markov chain condition $U\rightarrow X\rightarrow Y_{1},Y_{2}$, the more-capable condition implies that
@@
Note that the structure of the novel outer bound $\mathcal{R}_{o}$ given in Theorem 1, in particular, with respect to the constraints (9) and (10), is crucial for deriving the capacity results in Theorems 4 and 5. In the next section, we show that the bounds can also be used to derive capacity results for certain classes of Gaussian BCs.
@@ IV-A Converse
The set $\mathcal{R}_{o}^{G}$ constitutes an outer bound on the capacity region of the BC.
@@
Note that in the case of $\lambda=1$ (with $a=b$), or $\lambda=-1$ (with $a=-b$), the above equality is to be interpreted in the limiting sense. Further, when $\lambda=\frac{b}{a}$, the BC is physically degraded, for which the outer bound $\mathcal{R}_{o}^{G}$ is achievable. (See Remark 11.)
We now discuss the BC with perfectly correlated noises, i.e., $\lambda=1$ or $-1$, but with the assumption that $\lambda\neq\frac{b}{a}$.
@@ IV-B Capacity Regions for Specific Channels
$I(\hat{Y}_{2};Y_{2}|W,U,Y_{1})=\frac{1}{2}\psi\left(\frac{1-\lambda^{2}}{\hat{\sigma}^{2}}\right)=0$ (95)
@@
$I(X;Y_{1},\hat{Y}_{2})$ $=\frac{1}{2}\psi\left(\left(\frac{a^{2}+b^{2}-2\lambda ab+\hat{\sigma}^{2}a^{2}}{1-\lambda^{2}+\hat{\sigma}^{2}}\right)P\right)$
$=\frac{1}{2}\psi\left(\left(\frac{a^{2}+b^{2}-2\lambda ab+\hat{\sigma}^{2}a^{2}}{\hat{\sigma}^{2}}\right)P\right)$ (96)
@@
$I(X;Y_{1},\hat{Y_{2}}|W)$ $=\frac{1}{2}\psi\left(\beta\left(\frac{a^{2}+b^{2}-2\lambda ab+\hat{\sigma}^{2}a^{2}}{1-\lambda^{2}+\hat{\sigma}^{2}}\right)P\right)$
$=\frac{1}{2}\psi\left(\beta\left(\frac{a^{2}+b^{2}-2\lambda ab+\hat{\sigma}^{2}a^{2}}{\hat{\sigma}^{2}}\right)P\right).$ (97)
Note that if $|\lambda|=1$ but $\lambda\neq\frac{b}{a}$, then $a^{2}+b^{2}-2\lambda ab+\hat{\sigma}^{2}a^{2}\neq 0$. Therefore, by letting $\hat{\sigma}^{2}\rightarrow 0$, one can make the above two mutual information terms arbitrarily large, so they are would not be in effect in the characterization of the achievable rate region. For the case of $|a|\geq|b|$, the converse proof is readily given by $\mathcal{R}_{o}^{G}$ in Theorem 6. For the case of $|a|<|b|$, the achievable rate region (92)-(94) is maximized for $\beta=0$ and it coincides with the cut-set bound. ∎
@@
where $Z$ is a zero mean unit variance Gaussian noise. In this scenario, the two receivers see exactly the same noise but with a different sign. For this channel, the capacity result of Theorem 7 reduces to the following:
@@
Observe that in the case of common message only, even as $P\rightarrow 0$, it is possible to achieve a strictly positive rate of $R_{0}=\min\{C_{12},C_{21}\}$. This makes concrete the statement made in the introduction, namely, that it is possible to use an infinitesimally small amount of power to transmit information at a strictly positive rate. Note that the argument presented in the proof of Theorem 7 is only valid for strictly positive values of input power $P$. We summarize this observation in the following remark.
@@
For the special case of primitive relay channel with perfectly correlated noise, where $Y_{2}$ acts as a relay for $Y_{1}$, $C_{12}=0$ and there is only a private message for the first receiver, the capacity result given in Theorem 7 is reduced to $R_{1}\leq\psi\left(a^{2}P\right)+C_{21}$. In this case, using the quantize-bin-and-forward strategy, one can achieve the cut-set bound. Interestingly, the capacity does not depend on $b$ at all.
This result is an example of a semi-deterministic primitive relay channel, because the relay observation $Y_{2}=bX+Z_{2}$ is a deterministic function of the input $X$ and the receiver observation $Y_{1}=aX+Z_{1}$, when $Z_{1}$ and $Z_{2}$ are perfectly correlated and $\lambda\neq\frac{b}{a}$. In this case, the cut-set bound is achievable [32].
@@
where $Z$ is a zero mean unit variance Gaussian noise. In this case, the receiver $Y_{2}$ does not receive any information from the transmitter and only observes the additive noise of the receiver $Y_{1}$. For this channel, the capacity region can be derived by setting $a=1$ and $b=0$ in (92)-(94) and is given by
@@
In this case, the capacity region is given by
@@
This is the Gaussian counterpart of Example 2. Similar to Example 2, the receiver $Y_{1}$ (which needs to detect both common and private messages) does not receive any information from the transmitter directly. The capacity achieving cooperation protocol is that the receiver $Y_{2}$ first sends a compressed version of its received signal to the receiver $Y_{1}$ through the digital link $C_{21}$. Next, the receiver $Y_{1}$ decodes both messages using the information received (and its own signal, which is in fact just the channel noise) then forwards the bin index of the decoded common message to the receiver $Y_{2}$ through the link $C_{12}$. Lastly, the receiver $Y_{2}$ decodes the common message using the bin index received and its received signal.
@@
Theorem 7 is a capacity result for the Gaussian BC with degraded message sets, and with bidirectional conferencing decoders and perfectly correlated noises. Next, we establish the capacity region of a Gaussian BC with both common and private messages, and with one-sided cooperation.
@@
The proof of Theorem 8 is similar to that of Theorem 7 and therefore omitted.
@@
+$\qquad\qquad\qquad+\{C_{21}-1/2\}^{+}+C_{12}$ (114)

### v4

@@ Abstract
The two-user broadcast channel (BC) with receivers connected by bidirectional cooperation links of finite capacities, known as conferencing decoders, is considered. A novel capacity region outer bound is established based on multiple applications of the Csiszár-Körner identity. Achievable rate regions are derived by using Marton’s coding as the transmission scheme, together with different combinations of decode-and-forward and quantize-bin-and-forward strategies at the receivers. It is shown that the outer bound coincides with the achievable rate region for a new class of semi-deterministic BCs with degraded message sets; for this class of channels, one-round cooperation is sufficient to achieve the capacity. Capacity result is also derived for a class of more capable semi-deterministic BCs with both common and private messages and one-sided conferencing. For the Gaussian BC with conferencing decoders, if the noises at the decoders are perfectly correlated (i.e., the correlation is either 1 or -1), the new outer bound yields exact capacity region for two cases: i) BC with degraded message sets; ii) BC with one-sided conferencing from the weaker receiver to the stronger receiver. When the noises have arbitrary correlation, the outer bound is shown to be within half a bit from the capacity region for these same two cases. Finally, for the general Gaussian BC, a one-sided cooperation scheme based on decode-and-forward from the stronger receiver to the weaker receiver is shown to achieve the capacity region to within $\frac{1}{2}\log(\frac{2}{1-|\lambda|})$ bits, where $\lambda$ is the noise correlation. An interesting implication of these results is that for a Gaussian BC with perfectly negatively correlated noises and conferencing decoders with finite cooperation link capacities, it is possible to achieve a strictly positive rate using only an infinitesimal amount of transmit power.
@@
Broadcast channel, capacity region, correlated noise, primitive relay channel, receiver cooperation
@@ I-A Motivation
The situation is completely different if the receiver noises are correlated. Specifically, if the receiver noises are perfectly positively correlated, then cooperation would gain nothing, because $Y_{1}=Y_{2}$ in this case; as each receiver already knows what the other receiver knows, exchanging information does not help. However, if the receiver noises are perfectly negatively correlated, then cooperation can help significantly. Consider the case where $Y_{1}=X+Z$ and $Y_{2}=X-Z$, and the conferencing links have infinite capacities $C_{12}=C_{21}=\infty$. Such a BC with full receiver cooperation has infinite capacity at any positive transmit power $P$, because having access to both $Y_{1}$ and $Y_{2}$ allows the noises to be cancelled completely.
What happens if the noises are perfectly negatively correlated, but the conferencing links have finite capacity, e.g., $C_{12}=C_{21}=1$ bit per channel use? Complete noise cancellation is no longer possible. The analysis becomes more delicate, and this is one of the main topics of this paper. It turns out that an interesting phenomenon emerges—it is possible to achieve a strictly positive rate (i.e., 1 bit in the example above), even with an infinitesimally small amount of transmit power. In contrast to the single-user Gaussian channel where $P\to 0$ implies $C\to 0$, for this Gaussian BC with perfectly negatively correlated noises and conferencing decoders of fixed cooperative link capacities, we can have $P\to 0$, while $C\to 1$. In other words, the energy efficiency for transmitting one bit of information can be infinite!
Figure 1: The capacity of a Gaussian channel (left) goes to zero if the transmit power $P$ goes to zero at a fixed noise variance $N$. In a Gaussian BC with perfectly negatively correlated noises at the receiver (right), conferencing links of fixed capacities $C_{12}$ and $C_{21}$ between the receivers can allow a strictly positive rate to be achieved even at an infinitesimally small amount of transmit power.
@@ I-C Main Contributions
We then propose achievability strategies for the BC with conferencing decoders. In [20, Theorem 2] an achievable region is derived for the BC with bidirectional cooperation by applying Marton’s coding at the transmitter and the compress-and-forward cooperative scheme at both receivers. Another achievable region is given in [21] based on coding strategies for the partially cooperative relay broadcast channels. A third achievable region is given in [23, Appendix B] for the BC with one-sided cooperation between the receivers, which is derived by applying Marton’s coding at the transmitter and decode-and-forward as the cooperative protocol. All of these achievable schemes are however insufficient to either derive new capacity results for the discrete memoryless BCs, or approximate capacity results for the Gaussian BC.
@@ II Channel Model
For the BC with conferencing decoders, a length-$n$ code with $\Gamma$ conferencing rounds is described as follows. The transmitter encodes independent messages $M_{0}$, $M_{1}$, and $M_{2}$, which are uniformly distributed over the sets $1:2^{nR_{0}}]$, $1:2^{nR_{1}}]$, and $1:2^{nR_{2}}]$, respectively, into a codeword and sends over the channel according to the following:
@@ III-A Converse
for some joint PDFs $P(u,v,x)$, where $U,V\rightarrow X\rightarrow Y_{1},Y_{2}$ forms a Markov chain and $|\mathcal{U}|\leq|\mathcal{X}|+2$ and $|\mathcal{V}|\leq|\mathcal{X}|+2$. Then $\mathcal{R}_{o}$ constitutes an outer bound on the capacity region.
@@ III-B Achievability
$\zeta_{1}$ $=\{\alpha_{1}C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2})\}^{+}$ (17)
$\zeta_{2}$ $=\{C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})\}^{+}$ (18)
for some joint PDFs $P(u,v,w,x)P(\hat{y}_{1}|u,w,y_{1})P(\hat{y}_{2}|y_{2})$ and $\alpha_{1}\in[0,1]$. The convex closure of $\mathcal{R}_{i}^{(1)}$ is achievable.
@@
The achievable rate region $\mathcal{R}_{i}^{(1)}$ given in Theorem 2 clearly includes both of the regions previously presented in [20, Theorem 2] and [23, Appendix B] as subsets. Moreover, after some algebraic computation, one can show that $\mathcal{R}_{i}^{(1)}$, restricted over the distributions of the form $P(u,v,w,x)P(\hat{y}_{1}|w,y_{1})P(\hat{y}_{2}|y_{2})$, includes the achievable rate region of [21, Theorem 3] as a subset. Therefore, $\mathcal{R}_{i}^{(1)}$ encompasses these previously results. The achievable rate region $\mathcal{R}_{i}^{(1)}$ is evaluated over a larger set of distributions, i.e., $P(u,v,w,x)P(\hat{y}_{1}|u,w,y_{1})P(\hat{y}_{2}|y_{2})$, as compared to the one in [21, Theorem 3]. This improvement arises from the novel cooperation scheme at the first receiver, where it transmits a compressed version of the decoded signals U and W along with its received signal $Y_{1}$ to the second receiver. As discussed in [30], this novel approach may lead to derivation of new capacity results for certain types of Gaussian channels.
@@
Next, we present a second achievable rate region based on a combination of decode-and-forward and quantize-bin-and-forward at one receiver first and then quantize-bin-and-forward at the other receiver as the cooperation strategy. This strategy may be advantageous when part of the message is already decodable at the first receiver before conferencing.
@@
$\eta_{1}$ $=\{C_{12}-I(\hat{Y}_{1};U,Y_{1}|V,W,Y_{2})\}^{+}$ (40)
$\eta_{2}$ $=\{\alpha_{2}C_{21}-I(\hat{Y}_{2};Y_{2}|U,W,Y_{1})\}^{+}$ (41)
for some joint PDFs $P(u,v,w,x)P(\hat{y}_{1}|u,w,y_{1})P(\hat{y}_{2}|w,y_{2})$ and $\alpha_{2}\in[0,1]$. The convex closure of $\mathcal{R}_{i}^{(2)}$ is achievable.
@@
The cooperation strategy to achieve the rate region $\mathcal{R}_{i}^{(2)}$ is to apply Marton’s coding with a combination of decode-and-forward and quantize-bin-and-forward at the receiver $Y_{2}$ first, then quantize-bin-and-forward at the receiver $Y_{1}$. At the receiver $Y_{2}$, the capacity of the conferencing link $C_{21}$ is divided into two parts. A fraction $\bar{\alpha}_{2}C_{21}$ is devoted for decode-and-forward and the remaining $\alpha_{2}C_{21}$ is devoted for quantize-bin-and-forward scheme. The details of the proof are in Appendix C. ∎
@@
One can obtain an alternative achievable rate region by exchanging the cooperative protocol at the receivers in the achievability scheme of Proposition 2, i.e., to apply a combination of decode-and-forward and quantize-bin-and-forward at the receiver $Y_{1}$ first and then quantize-bin-and-forward at the receiver $Y_{2}$.
@@ III-C Capacity Regions for Specific Channels
The achievability is derived by setting $U=X$, $W=V$, $\hat{Y}_{1}=\varnothing$, and $\hat{Y}_{2}=Y_{2}$ in $\mathcal{R}_{i}^{(1)}$. The converse proof is readily given by $\mathcal{R}_{o}$. ∎
@@
Figure 3: Capacity region and cut-set bound for the BC in Example 2, where $Z$ is a binary noise with P(Z=0)=0.2{P(Z=0)=0.2}.
Unlike the previous example, for which the capacity region coincides with the cut-set bound, the capacity region for this BC is characterized using an auxiliary variable $V$. In Fig. 3, we plot the capacity region along with the cut-set bound for the channel in Example 2 with $P(Z=0)=0.2$ for two cases: $(C_{12},C_{21})=(0,0.9)$ and $(C_{12},C_{21})=(0.1,0.9)$, by searching over all possible $P(v,x)$. As can be seen, the capacity region lies strictly below the cut-set bound in this example. Moreover, $C_{12}$ strictly improves the capacity region, although $Y_{1}$ only observes the noise. In fact, the constraint (69) (which arises from (60)) plays a crucial role in characterizing the capacity region. Without (69), the remaining constraints simply reduce to the cut-set bound.
For this channel, the optimal coding strategy is as follows. Note that the receiver $Y_{1}$ (which is supposed to decode both common and private messages) does not receive any information from the transmitter directly. The optimal cooperation strategy is that the receiver $Y_{2}$ first sends a compressed version of its received signal to the receiver $Y_{1}$ through the digital link $C_{21}$. Next, the receiver $Y_{1}$ decodes both messages using the information received (and its own signal which is in fact the channel noise) and then performs decode-and-forward of the common message through the link $C_{12}$ to the receiver $Y_{2}$ (using the variable $V$). Lastly, the receiver $Y_{2}$ decodes the common message using also its own signal.
It is worthwhile to make the following observation. In Example 2, as the receiver $Y_{1}$ observes the channel noise only, one might think that a cooperative scheme in which $Y_{1}$ applies quantize-bin-and-forward and $Y_{2}$ applies decode-and-forward is the right strategy. However, it turns out that such a scheme is not optimal when the message to $Y_{2}$ is a degraded version of the message to $Y_{1}$. This is in contrast to Example 1, in which the message set degrades in the opposite direction and performing quantize-bin-and-forward at the receiver that observes the noise only is the capacity-achieving strategy.
As far as we know, the above results are the first cases where a combination of quantize-bin-and-forward and decode-and-forward strategies yields an optimal bidirectional cooperation protocol. There is no previously known capacity result in the literature for channels with bidirectional cooperation between the receivers. Moreover, these results demonstrate that a one-round cooperation scheme is sufficient for achieving capacity for this class of channels.
@@
The achievability is derived by setting $U=X$, $W=V$, $\hat{Y}_{1}=\varnothing$, and $\hat{Y}_{2}=Y_{2}$ in $\mathcal{R}_{i}^{(1)}$. Note that the more capable condition and one-sided conferencing imply that the private message for user 2 can always be decoded by user 1.
For the converse, note that due to the Markov chain condition $U\rightarrow X\rightarrow Y_{1},Y_{2}$, the more-capable condition implies
@@
Remark 7.
The structure of the novel outer bound $\mathcal{R}_{o}$ given in Theorem 1, in particular, with respect to the constraints (9) and (10), is crucial for deriving the capacity results in Theorems 4 and 5. Specific combinations of conditions, such as semi-determinism, degraded message sets, and the more capable channel, make it possible for the converse to be tight.
@@ IV-A Converse
Then $\mathcal{R}_{o}^{G}$ constitutes an outer bound on the capacity region of the BC.
@@
Note that in the case of $\lambda=1$ (with $a=b$), or $\lambda=-1$ (with a=ba=-b), the above equality is to be interpreted in the limiting sense. Further, when $\lambda=\frac{b}{a}$, the BC is physically degraded, for which the outer bound $\mathcal{R}_{o}^{G}$ is achievable. (See Remark 12.)
We now discuss the BC with perfectly correlated noises, i.e., $\lambda=1$ or 1-1, but with the assumption that $\lambda\neq\frac{b}{a}$.
@@ IV-B Capacity Regions for Specific Channels
$I(\hat{Y}_{2};Y_{2}|W,U,Y_{1})=\psi\left(\frac{1-\lambda^{2}}{\hat{\sigma}^{2}}\right)=0$ (95)
@@
$I(X;Y_{1},\hat{Y}_{2})$ $=\psi\left(\left(\frac{a^{2}+b^{2}-2\lambda ab+\hat{\sigma}^{2}a^{2}}{1-\lambda^{2}+\hat{\sigma}^{2}}\right)P\right)$
$=\psi\left(\left(\frac{a^{2}+b^{2}-2\lambda ab+\hat{\sigma}^{2}a^{2}}{\hat{\sigma}^{2}}\right)P\right)$ (96)
@@
$I(X;Y_{1},\hat{Y_{2}}|W)$ $=\psi\left(\beta\left(\frac{a^{2}+b^{2}-2\lambda ab+\hat{\sigma}^{2}a^{2}}{1-\lambda^{2}+\hat{\sigma}^{2}}\right)P\right)$
$=\psi\left(\beta\left(\frac{a^{2}+b^{2}-2\lambda ab+\hat{\sigma}^{2}a^{2}}{\hat{\sigma}^{2}}\right)P\right).$ (97)
Note that if $|\lambda|=1$ but $\lambda\neq\frac{b}{a}$, then $a^{2}+b^{2}-2\lambda ab+\hat{\sigma}^{2}a^{2}\neq 0$. Therefore, by letting $\hat{\sigma}^{2}\rightarrow 0$, one can make the above two mutual information terms arbitrarily large, so they would not be in effect in the characterization of the achievable rate region. For the case of $|a|\geq|b|$, the converse proof is readily given by $\mathcal{R}_{o}^{G}$ in Theorem 6. For the case of $|a|<|b|$, the achievable rate region (92)-(94) is maximized for $\beta=0$ and it coincides with the cut-set bound. ∎
@@
where $Z$ is a zero-mean unit-variance Gaussian noise. In this scenario, the two receivers see exactly the same noise but with a different sign. The capacity region of this channel with degraded message sets can be obtained from Theorem 7 as:
@@
Observe that in the case of common message only, even as $P\rightarrow 0$, it is possible to achieve a strictly positive rate of $R_{0}=\min\{C_{12},C_{21}\}$. This makes concrete the statement made in the introduction of this paper, namely, that it is possible to use an infinitesimally small amount of power to transmit information at a strictly positive rate. Note that the argument presented in the proof of Theorem 7 is only valid for strictly positive values of input power $P$. We summarize this observation in the following remark.
@@
For the special case of a primitive relay channel with perfectly correlated noises, i.e., $|\lambda|=1$, where $Y_{2}$ acts as a relay for $Y_{1}$, $C_{12}=0$, and there is only a private message for the first receiver, the capacity result given in Theorem 7 is reduced to $R_{1}\leq\psi\left(a^{2}P\right)+C_{21}$. In this case, using the quantize-bin-and-forward strategy, one can achieve the cut-set bound. Interestingly, the capacity does not depend on $b$ at all (as long as $\lambda\neq\frac{b}{a}$).
This result is an example of a semi-deterministic primitive relay channel, because the relay observation $Y_{2}=bX+Z_{2}$ is a deterministic function of the input $X$ and the receiver observation $Y_{1}=aX+Z_{1}$, when $Z_{1}$ and $Z_{2}$ are perfectly correlated (i.e., $|\lambda|=1$) and $\lambda\neq\frac{b}{a}$. In this case, the cut-set bound is achievable [33].
@@
where $Z$ is a zero-mean unit-variance Gaussian noise. In this case, the receiver $Y_{2}$ does not receive any information from the transmitter and only observes the additive noise of the receiver $Y_{1}$. The capacity region of this channel with degraded message sets can be derived by setting $a=1$, $b=0$ in (92)-(94) as:
@@
The capacity region of this channel with degraded message sets is given by
@@
This is the Gaussian counterpart of Example 2. Similar to Example 2, the receiver $Y_{1}$ (which needs to detect both common and private messages) does not receive any information from the transmitter directly. The capacity achieving cooperation protocol is that the receiver $Y_{2}$ first sends a compressed version of its received signal to the receiver $Y_{1}$ through the digital link $C_{21}$. Next, the receiver $Y_{1}$ decodes both messages using the information received (and its own signal, which is in fact just the channel noise) then forwards the bin index of the decoded common message to the receiver $Y_{2}$ through the link $C_{12}$. Lastly, the receiver $Y_{2}$ decodes the common message using the received bin index and its own received signal.
@@
Theorem 7 is a capacity result for the Gaussian BC with degraded message sets, and with bidirectional conferencing decoders and perfectly correlated noises. Next, we treat the case of a Gaussian BC with both common and private messages, but with one-sided cooperation.
@@
Proof:
The proof is similar to that of Theorem 7. ∎
@@
$\qquad\qquad\qquad+\{C_{21}-1/2\}^{+}+C_{12}$ (114)